OpenStudy (anonymous):

ALG II Ashley solved the exponential equation 3x+1 = 15 and her work is shown below. What is the first step she did incorrectly? Step 1: log 3x+1 = log15 Step 2: (x + 1)log 3 = log15 Step 3: x + 1 = log 15/log 3 Step 4: x+1 = 2.708050/1.098612 Step 5: x + 1 = 1.418858 Step 6: x = 0.418858

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle3^{x+1}=15 }\) like this?

OpenStudy (solomonzelman):

That is not your original equation, is it?

OpenStudy (anonymous):

It is.. I must've forgotten the square root arrow.

OpenStudy (solomonzelman):

Wait, what is your original question then?

OpenStudy (anonymous):

It's already solved, but I have to find out which step is incorrect.

OpenStudy (solomonzelman):

What is your initial equation that you're solving?

OpenStudy (solomonzelman):

If I don't know the initial question, then I can't help you...

OpenStudy (anonymous):

\[3^{^{x+1}}=15\]

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle3^{x+1}=15 }\) \(\large\color{black}{ \displaystyle3^{x+1}\div 3^{1}=15\div 3^{1} }\) \(\large\color{black}{ \displaystyle3^{x}=5 }\) \(\large\color{black}{ \displaystyle\log_3(3^{x})=\log_3(5) }\) \(\large\color{black}{ \displaystyle x\log_3(3)=\log_3(5) }\) \(\large\color{black}{ \displaystyle x=\log_3(5)\approx 1.4649735207 }\)

OpenStudy (solomonzelman):

this is one way to do it, but lets check yours...

OpenStudy (solomonzelman):

Step 1: log 3x+1 = log15 Step 2: (x + 1)log 3 = log15 Step 3: x + 1 = log 15/log 3 \(\color{red}{\rm Step~4:}\) x + 1 = log (5•3)/log (3) \(\color{red}{\rm Step~5:}\) x + 1 = [ log (5) +log(3) ]/log (3) \(\color{red}{\rm Step~6:}\) x + 1 = log (5)/log (3) + log (3)/log (3) \(\color{red}{\rm Step~7:}\) x + 1 = log (5)/log (3) + 1 \(\color{red}{\rm Step~8:}\) x = log (5)/log (3) \(\color{red}{\rm Step~9:}\) x = 1.4649735207

OpenStudy (anonymous):

She went wrong at Step 4?

OpenStudy (solomonzelman):

wherever the error was, the error is clearly wrong.

OpenStudy (solomonzelman):

The error is that: log(3)=0.477

OpenStudy (solomonzelman):

So yes, Step 4 is wrong.